FFC operational status:
IDA engineers visited the station and removed the STS-1 seismometers and reconfigured the T360SV to the FFC00 Q330. A Trillium240 was installed as the secondary sensor on the FFC10 Q330.
A Nanometrics engineer visited the station and installed a T360SV seismometer.
The station is back online day 2020,246 after the station operator replaced the LAN extenders between the vault and recording room.
The station is down as of day 2020,232 due to communication issues between the vault and recording room. IDA engineers are working with the station operator to resolve these issues.
The station is back online after the station operator replaced the ethernet extenders and surge protectors.
The station is down as of day 2016,214-04:04 due to communication problems between the vault and recording room. The station operator reports there was an electrical storm in the area.
The station is back online as of day 2016,180-13:17. There were no known actions taken at the station, however the station operator noted a lightning storm in the area.
The station is down as of day 2016,114-15:43 due to communication problems between the station computer and digitizers. The station operator has been requested to investigate.
The station operator replaced the ffc10 digitizer with one shipped from UCSD. Short gaps have ceased as of day 2016,062-16:12.
There are short gaps in all channels from the ffc10 digitizer. After swapping cables from the digitizers to the ethernet switch the gapping remained indicating a faulty digitizer. A new digitizer is being prepared to ship to the station.
The station is back online as of day 2016,025-15:56 after the station operator replaced the ethernet switch and LAN extenders shipped from UCSD.
The station is down as of day 2015,348-21:09 due to a communication problem between the digitizers and the station computers.
An IDA engineer visited the station and replaced the Trillium240 seismometer and cable. The north-south component of the Trillium 240 seismometer had an anomalous diminished output beginning around 2013,359. The sensor was taken for testing and a temporary replacement was installed in September 2014. Tests of the original sensor did not show any definitive problems but the seismometer’s mixer board was replaced. Upon re-installation with the existing cable the output of the sensor decreased about 50%. When the cable was switched with a new one sent back with the original sensor the output improved.
An IDA engineer is on-site to replace one of the Q330HR digitizers and the Trillium 240 seismometer.
Telemetry to the station has been restored.
Telemetry to the station is down since 162-06:09. The station operator has been requested to investigate.
There are communication problems between the digitizers and station computer causing small gaps in the data since day 2013,091-16:55.
Telemetry has been restored. A technician from the service provider also visited the station and adjusted the satellite dish and verified signal levels.
Telemetry to the station is down as of day 2012,211-23:22. A faulty satellite modem is suspected. IDA engineers are working with the station operator to resolve the issue.
The station is back online as of day 2011,138-13:19 after the station operator replaced the power supply IDA engineers shipped from UCSD.
Telemetry to the station was down on day 2011,116-04:01 due to a power outage. The station operator checked on the station and data acquisition resumed at 2011,116-21:22. However, power problems persist. The station is down.
The station is back on-line as of day 2011,042-21:10.
The station is down as of day 2011,040-07:26 due to communication problems between the DAS and the ISP. IDA engineers are working with the station operator to troubleshoot.
The vertical STS-1 sensor has a vacuum seal failure. The station operator has evacuated the sensor twice but the vacuum was lost over time.
The station operator swapped the two STS-1 horizontal sensor cables back to their original configuration at about day 2009,336-21:51. The station operator also evacuated the vertical STS-1 sensor but the seal is not holding.
There is excessive noise on the vertical and one of the horizontal channels from the STS-1 seismometers. To investigate the source of this noise the station operator switched the cables to the two STS-1 horizontal sensors on day 2009,317-22:07. The QC’d data has been corrected for the switched cables accordingly.
The station has been upgraded to a Quanterra Q330HR data acquisition system. Sensors installed include STS-1 broadband and Trillium 240 short period seismometers, and a strong motion sensor.
The station is back on-line as of about day 2009,132 after the station operator cycled power to the DAS and the ARS.
The station is down due to problems with the ARS that are yet to be determined. The station operator has been requested to investigate.
The station is back up as of day 2007,229 after an IDA engineer visited the station and replaced the ISP.
The station is down due to problems with the ISP that are yet to be determined.
The station was down from day 273 (Sept.30) to 285 (Oct.12) due to problems with the DAS and the telemetry link to the station. Data acquisition resumed when telemetry was restored and the DAS was rebooted.
An IDA engineer visited the station to upgrade the ISP and improve telemetry.
Telemetry system upgrade to VSAT complete (Sept.10).
Station is up after DAS communication board problem was corrected by an IDA engineer.
Station is down due to problems with the DAS. “No Data from DAS” message continued after DAS and ARS reset.
Vertical channels from the STS-1 sensor are back to normal after the station operator checked and corrected the vacuum for this instrument.
Vertical channels from the STS-1 sensor have excessive noise, for reasons not yet determined.
Station operator reports installation of replacement modem and surge suppressor. Telemetry restored.
Telemetry to the station down after a severe storm, modem damaged.
New DAT drive shipped.
Operator reports he has replaced the damaged modem with the unit shipped to him. Telemetry has been restored.
Operator reports the modem has been destroyed during a power failure that effected the area’s telephone system.
New EPROMs installed to correct DSP problem.
Station is operating but recording stops from time to time. DAS reset remotely from San Diego on 2002.04.22.